OpenStudy (anonymous):

What is equal to 6√729x^7 assume x is non-negative

OpenStudy (anonymous):

\[6\sqrt{729} * x^7\] Is this it?

OpenStudy (anonymous):

\[\sqrt[6]{729}*x^7\] or this?

OpenStudy (anonymous):

@RyanL. ^6 √(729x^7)

OpenStudy (anonymous):

ok you can rewrite the expression knowing this \[\sqrt[a]{x^b}=x^\frac{b}{a}\]

OpenStudy (anonymous):

729x7/6 (7/6 is fraction no division)

OpenStudy (anonymous):

@RyanL.

OpenStudy (anonymous):

what do you mean by no division?

OpenStudy (anonymous):

not division. I dont know how to make a fraction on this computer.

OpenStudy (anonymous):

\[(729x^7)^\frac{1}{6}\]

OpenStudy (anonymous):

A: 2x ^6√3x B: 3x ^6√x C: x ^6√7x C: 3x^6√2x

OpenStudy (anonymous):

@RyanL.